### v1

# The Hessian of tall-skinny networks is easy to invert

Abstract
We describe an exact algorithm for solving linear systems $Hx=b$ where $H$ is the Hessian of a deep net. The method computes Hessian-inverse-vector products without storing the Hessian or its inverse in time and storage that scale linearly in the number of layers. Compared to the naive approach of first computing the Hessian, then solving the linear system, which takes storage that’s quadratic in the number of parameters and cubically many operations, our Hessian-inverse-vector product method scales roughly like Pearlmutter’s algorithm for computing Hessian-vector products.
1 Introduction
The Hessian of a deep net is the matrix of second-order mixed partial derivatives of its loss with respect to its parameters. Decades ago, when deep nets had only hundreds or thousands of parameters, the Hessian matrix could be inverted to implement optimizers that converged much faster than gradient descent [Watrous1987, Barnard1992]. But for large modern deep nets, using the Hessian has become less practical: The Hessian of a model with a billion parameters would have a quintillion entries, which is far larger than can be stored, multiplied, factorized, or inverted, even in the largest data centers. A common workaround is to approximate the Hessian as a low-rank matrix [Webb1988, lbfgs-deep-net] or as a diagonal matrix [becker-lecun-89, adagrad, adam]. Such approximations make it easier to apply the inverse of the Hessian to a vector. This article shows how to compute and apply the inverse of the Hessian exactly without storing the Hessian and without storing its inverse. The Hessian-inverse-vector product can be computed in time and storage that scales linearly with the number of layers in the model, and cubically in the number of parameters and activations in each layer.
Pearlmutter [pearlmutter94] showed the product of the Hessian with a fixed vector (the so-called Hessian-vector product) can be computed in time and storage that scales linearly with the number of layers in the network. The scaling behavior of Pearlmutter’s method is much better than the cubic scaling of the naive algorithm that first computes the Hessian matrix and then multiplies it by the vector. The method transforms the original network into a new network whose gradient is the desired Hessian-vector product. To compute the Hessian-vector product, one then just applies backpropagation to the new network. Given a way to compute the Hessian-vector product, one can indirectly compute the Hessian-inverse-vector product via, say Krylov iterations like Conjugate Gradient. However, the quality of the result would then depend on the conditioning of the Hessian, which is notoriously poor for deep nets [ying-behrooz-hessian-spectrum]. Unfortunately, there seems to exist no variant of Pearlmutter’s trick to compute the Hessian-inverse-vector products directly.
This article offers a way to compute Hessian-inverse-vector products using roughly as much computation and storage as Pearlmutter’s algorithm requires to compute the Hessian-vector product. The Hessian of a deep net has a regular structure that makes it possible to compute Hessian-inverse-vector products exactly and efficiently. Regardless of the specific operations in each layer, the Hessian of a layerwise deep net is a second order matrix polynomial that involves the first order and second order mixed derivatives of each layer, and the inverse of a block-bi-diagonal operator that represents the backpropagation algorithm. Exploiting this structure offers a direct way to compute the Hessian-vector product using exactly the same operations as Pearlmutter’s algorithm (see Appendix B). It also leads to a way to compute Hessian-inverse-vector product that does not require forming or storing the full Hessian. Assume an oracle that offers the first order and second order mixed derivatives of each layer. For an $L$-layer deep net where each layer has $p$ parameters and produces $a$ activations, naively storing the Hessian would require $O(L^{2}p^{2})$ memory, and solving a linear system would require $O\left(L^{3}p^{3}\right)$ operations in addition to the oracle queries. In contrast, we will show how to perform these operations using only $O(L\max(a,p)^{3})$ additional computation. The dependence on the parameters on the number of activations and parameters in each layer is still cubic, but the dependence on the number of layers is only linear. This makes operating on the Hessian of tall and skinny networks more efficient than the Hessian of short and fat networks.
Although modern networks are typically short and wide, our method runs faster than the naive hessian-inverse-vector product algorithm on tall and skinny networks, since it ameliorates the naive algorithm’s cubic dependence on depth to a mild linear dependence on depth. A follow-up article will explore training tall-skinny networks using the Hessian-inverse-vector product as a preconditioner. We expect that a significant training speedup will motivate a return to deeper network designs.
2 Overview
Our objective is to efficiently solve linear systems of the form $Hx=g$, where $H$ is the Hessian of a deep neural network, without forming $H$ explicitly. To do this, we employ the following strategy:
  1. 1.
    Write down the gradient of the deep net in matrix form, as a bi-diagonal system of linear equations. Solving this system uses back-substitution, which in this case coincides exactly with the computations carried out by backpropagation.
  2. 2.
    Differentiate this matrix form to obtain an expression for the Hessian. This expression involves a second-order polynomial in the inverse of the aforementioned bi-diagonal matrix.
  3. 3.
    Lift this polynomial into a higher-dimensional matrix form. After pivoting, this matrix becomes block-tri-diagonal.
  4. 4.
    Factorize the resulting block-tri-diagonal into a lower bi-diagonal block matrix, a diagonal block matrix and an upper block-diagonal matrix (the LDU factorization).
  5. 5.
    Solve the resulting system using forward and backward substitution. This is similar to running backpropagation on a modified version of the original network that’s designed to compute the Hessian-inverse-vector product.
  6. 6.
    Un-pivot the result, and report the un-lifted solution.
Algorithm 1 summarizes the steps above.
Algorithm 1 Compute the hessian-inverse-vector product by solving $(H+\epsilon I)x=b$
1:A vector $b$, damping constant $\epsilon$.
[…]
3:Define the sparse block matrix $\mathcal{K}$ and the augmented system as in Equation 13.
4:Pivot into permuted system $\mathcal{K}^{\prime}x^{\prime}=b^{\prime}$ where $\mathcal{K}^{\prime}=\Pi\mathcal{K}\Pi^{\top}$ is block-tri-diagonal following Equation (14).
5:Solve $\mathcal{K}^{\prime}x^{\prime}=b^{\prime}$ using block-LDU decomposition.
6:Recover the solution $x$ from the permuted vector $\Pi\left[\begin{smallmatrix}x\\ y\\ z\end{smallmatrix}^{\prime}\right]$.
The algorithm is straightforward to implement with a suitable library of block matrix operations. [^1] https://github.com/a-rahimi/hessian
3 Notation
We write a deep net as a pipeline of functions $\ell=1,\ldots,L$,
$z_{1}$ $=f_{1}(z_{0};x_{1})$
$\ldots$
$z_{\ell}$ $=f_{\ell}(z_{\ell-1};x_{\ell})$
$\ldots$
$z_{L}$ $=f_{L}(z_{L-1};x_{L})$ (1)
The vectors $x_{1},\ldots,x_{L}$ are the parameters of the pipeline. The vectors $z_{1},\ldots,z_{L}$ are its intermediate activations, and $z_{0}$ is the input to the pipeline. The last layer $f_{L}$ computes the final activations and their training loss, so the scalar $z_{L}$ is the loss of the model on the input $z_{0}$. To make this loss’s dependence on $z_{0}$ and the parameters explicit, we sometimes write it as $z_{L}(z_{0};x)$. This formalization deviates slightly from the traditional deep net formalism in two ways: First, the training labels are subsumed in $z_{0}$, and are propagated through the layers until they’re used in the loss. Second, the last layer fuses the loss (which has no parameters) and the last layer (which does).
We’ll assume the first and partial derivatives of each layer with respect to its parameters and its inputs exist. This poses some complications with ReLU activations and other non-differentiable operations in modern networks. Notably, for the Hessian to be symmetric, it must be differentiable everywhere. We’ll largely ignore this complication and assume that differentiable approximations to these operations are used.
At the end of each section, we’ll count the number of floating point operations required to compute various expressions. While the derivations do not impose any restrictions on the shape of the layers, for the purpose of this accounting, we’ll assume all but the last $L$ layers have $a$-dimensional activations ($z_{\ell}\in\mathop{\mathbb{R}}^{a}$) and $p$-dimensional parameters ($x_{\ell}\in\mathop{\mathbb{R}}^{p}$).
4 Backpropagation, the matrix way
We would like to fit the vector of parameters $x=(x_{1},\ldots,x_{L})$ given a training dataset, which we represent by a stochastic input $z_{0}$ to the pipeline. Training the model proceeds by gradient descent steps along the stochastic gradient $\partial z_{L}(z_{0};x)/\partial x$. The components of this direction can be computed by the chain rule with a backward recursion:
$\frac{\partial z_{L}}{\partial x_{\ell}}$ $=\underbrace{\frac{\partial z_{L}}{\partial z_{\ell}}}_{b_{\ell}}\underbrace{\frac{\partial z_{\ell}}{\partial x_{\ell}}}_{\nabla_{x}f_{\ell}}$ (2)
$\frac{\partial z_{L}}{\partial z_{\ell}}$ $=\underbrace{\frac{\partial z_{L}}{\partial z_{\ell+1}}}_{b_{\ell+1}}\underbrace{\frac{\partial z_{\ell+1}}{\partial z_{\ell}}}_{\nabla_{z}f_{\ell+1}}.$ (3)
The identification $b_{\ell}\equiv\frac{\partial z_{L}}{\partial z_{\ell}}$, $\nabla_{x}f_{\ell}\equiv\frac{\partial z_{\ell}}{\partial x_{\ell}}$, and $\nabla_{z}f_{\ell}\equiv\frac{\partial z_{\ell}}{\partial z_{\ell-1}}$ turns this recurrence into
$\frac{\partial z_{L}}{\partial x_{\ell}}$ $=b_{\ell}\cdot\nabla_{x}f_{\ell}$ (4)
$b_{\ell}$ $=b_{\ell+1}\cdot\nabla_{z}f_{\ell+1},$ (5)
with the base case $b_{L}=1$, a scalar. These two equations can be written in vector form as
$\frac{\partial z_{L}}{\partial x}=\begin{bmatrix}\frac{\partial z_{L}}{\partial x_{1}}&\cdots&\frac{\partial z_{L}}{\partial x_{L}}\end{bmatrix}=\underbrace{\begin{bmatrix}b_{1}&\cdots&b_{L}\end{bmatrix}}_{\equiv b}\underbrace{\begin{bmatrix}\nabla_{x}f_{1}&\\ &\ddots\\ &&\nabla_{x}f_{L}\end{bmatrix}}_{\equiv D_{x}},$ (6)
and
$\begin{bmatrix}b_{1}&b_{2}&b_{3}&\cdots&b_{L-1}&b_{L}\end{bmatrix}\underbrace{\begin{bmatrix}I\\ -\nabla_{z}f_{2}&I\\ &-\nabla_{z}f_{3}&I\\ &&\ddots&\ddots\\ &&&-\nabla_{z}f_{L}&1\\ \end{bmatrix}}_{\equiv M}=\underbrace{\begin{bmatrix}0&\cdots&1\end{bmatrix}}_{\equiv e_{L}}.$ (7)
Solving for $b$ and substituting back gives
$\frac{\partial z_{L}}{\partial x}=e_{L}M^{-1}D_{x}.$ (8)
The matrix $M$ is block bi-diagonal. Its diagonal entries are identity matrices, and its off-diagonal matrices are the gradient of the intermediate activations with respect to the layer’s parameters. The matrix $D_{x}$ is block diagonal, with the block as the derivative of each layer’s activations with respect to its inputs. $M$ is invertible because the spectrum of a triangular matrix can be read off its diagonal, which in this case is all ones.
5 The Hessian
To obtain the Hessian, we use similar techniques to compute the gradient of Equation (8) with respect to $x$. The gradient we computed in Equation (8) is the unique vector $g$ such that $dz_{L}\equiv z_{L}(x+dx)-z_{L}(dx)\to g(x)^{\top}dx$ as $dx\to 0$. Similarly, the Hessian $H$ of $z_{L}$ with respect to the parameters is the unique matrix $H(x)$ such that $dg\equiv g(x+dx)-g(x)\to H(x)\;dx$ as $dx\to 0$. Appendix A shows that the Hessian has the following form:
Claim 1.
The Hessian of the loss $z_{L}$ with respect to the vector of parameters $x$ is given by
$H=D_{D}D_{xx}+D_{D}D_{zx}PM^{-1}D_{x}+D_{x}^{\top}M^{-\top}P^{\top}D_{M}D_{xz}+D_{x}^{\top}M^{-\top}P^{\top}D_{M}D_{zz}PM^{-1}D_{x},$ (9)
with the following matrices:
$D_{D}$ $\equiv\begin{bmatrix}\underbrace{I\otimes b_{1}}_{p\times ap}&&\\ &\ddots&\\ &&I\otimes b_{L}\end{bmatrix},\qquad D_{M}\equiv\begin{bmatrix}\underbrace{I\otimes b_{1}}_{a\times a^{2}}&&\\ &\ddots&\\ &&I\otimes b_{L}\end{bmatrix}$,
$P$ $\equiv\begin{bmatrix}0\\ I&0\\ &\ddots\\ &I&0\end{bmatrix},\qquad M\equiv\begin{bmatrix}I\\ \underbrace{-\nabla_{z}f_{2}}_{a\times a}&I\\ &-\nabla_{z}f_{3}&I\\ &&\ddots&\ddots\\ &&&-\nabla_{z}f_{L}&1\\ \end{bmatrix}$,
$D_{x}$ $\equiv\begin{bmatrix}\underbrace{\nabla_{x}f_{1}}_{a\times p}&\\ &\ddots\\ &&\nabla_{x}f_{L}\end{bmatrix},\qquad D_{xx}\equiv\begin{bmatrix}\underbrace{\nabla_{xx}f_{1}}_{ap\times p}&&\\ &\ddots&\\ &&\nabla_{xx}f_{L}\end{bmatrix},\qquad D_{xz}\equiv\begin{bmatrix}\underbrace{\nabla_{xz}f_{1}}_{a^{2}\times p}&&\\ &\ddots&\\ &&\nabla_{xz}f_{L}\end{bmatrix}$,
$D_{zx}$ $\equiv\begin{bmatrix}\underbrace{\nabla_{zx}f_{1}}_{ap\times a}&&\\ &\ddots&\\ &&\nabla_{zx}f_{L}\end{bmatrix},\qquad D_{zz}\equiv\begin{bmatrix}\underbrace{\nabla_{zz}f_{1}}_{a^{2}\times a}&&\\ &\ddots&\\ &&\nabla_{zz}f_{L}\end{bmatrix}$.
Given a vector $x\in\mathop{\mathbb{R}}^{Lp}$, the formula above allows us to compute $Hx$ in $O\left(Lap^{2}+La^{2}p+La^{3}\right)$ operations without forming $H$. This cost is dominated by multiplying by the $D_{xx}$, $D_{zx}$, and $D_{zz}$ matrices. Appendix B shows that these operations are exactly the operations performed in Pearlmutter’s trick to compute the Hessian-vector product.
To solve systems of the form $Hx=b$, one could use Krylov methods to repeatedly multiply by $H$ without forming $H$ (a possibility Pearlmutter considered [pearlmutter94]). This would require applying $H$ some number of times that depends on the condition number of $H$. However, the next section shows how to solve systems of the form $H^{-1}x=b$ with only $\max(a,p)$ times more operations than are needed to compute $Hx$.
6 Applying the inverse of the Hessian
The above shows that the Hessian is a second order matrix polynomial in $M^{-1}$. While $M$ itself is block-bidiagonal, $M^{-1}$ is dense, so $H$ is dense. Nevertheless, this polynomial can be lifted into a higher order object whose inverse is easy to compute:
$H=D_{D}D_{xx}+D_{D}D_{zx}PM^{-1}D_{x}+D_{x}^{\top}M^{-\top}P^{\top}D_{M}D_{xz}+D_{x}^{\top}M^{-\top}P^{\top}D_{M}D_{zz}PM^{-1}D_{x}.$
In general, $H$ is singular. We wish to solve the system $(H+\epsilon I)x=g$ for $x$. We can convert this dense system involving inverses into a larger, sparse system by introducing auxiliary variables. Define
$y\equiv M^{-1}D_{x}x,$ (10)
which implies $My-D_{x}x=0$. Define a second auxiliary variable
$z\equiv M^{-\top}\left(P^{\top}D_{M}D_{xz}x+P^{\top}D_{M}D_{zz}Py\right),$ (11)
which implies $M^{\top}z-P^{\top}D_{M}D_{xz}x-P^{\top}D_{M}D_{zz}Py=0$. With these substitutions, $(H+\epsilon I)x=g$ becomes
$\left(D_{D}D_{xx}+\epsilon I\right)x+D_{D}D_{zx}Py+D_{x}^{\top}z=g.$ (12)
Collecting these three linear equations gives us a unified system
$\underbrace{\begin{bmatrix}D_{D}D_{xx}+\epsilon I&D_{D}D_{zx}P&D_{x}^{\top}\\ -D_{x}&M&0\\ -P^{\top}D_{M}D_{xz}&-P^{\top}D_{M}D_{zz}P&M^{\top}\end{bmatrix}}_{\mathcal{K}}\begin{bmatrix}x\\ y\\ z\end{bmatrix}=\begin{bmatrix}g\\ 0\\ 0\end{bmatrix}$. (13)
Finding $x=\left(H+\epsilon I\right)^{-1}g$ is equivalent to solving this block-linear system and reporting the resulting $x$. The benefit of doing this is that this system can be pivoted into a block-tri-diagonal system, which can be solved more efficiently than Gaussian elimination on $\mathcal{K}$ or $H$.
The pivoting we’ll apply reorders the blocks of the variables $x,y,z$ from $x_{1},\ldots,x_{L},y_{1},\ldots,y_{L},z_{1},\ldots,z_{L}$ to $x_{1},y_{1},z_{1},\dots,x_{L},y_{L},z_{L}$. This pivoting operation acts as a kind of transpose operation on block matrices (its generalization is called a “commutation matrix” in [magnus-neudecker]). When applied to $\mathcal{K}$, it turns it into a block-tri-diagonal matrix. To get a better feel for this operation, denote the $ij$th block of $\mathcal{K}$ by $\mathcal{K}_{ij}$ and the $uv$th sub-block of this block by $\mathcal{K}_{ij,uv}$. Similarly, $x$ denote a vector conformant with $\mathcal{K}$, with $x_{j}$ denoting the column vector that multiplies each block $\mathcal{K}_{\cdot j}$ and denote the $v$th sub-block of $x_{j}$ by $x_{jv}$, which multiplies $\mathcal{K}_{\cdot j,\cdot v}$. We say that $x$ is $j$-major because as we traverse the entries of $x$ from top to bottom, the index $j$ increments more slowly than the $v$ index.
Denote by $\Pi$ the permutation that transposes the blocks with the sub-blocks. Applying this permutation matrix $\Pi$ to $x$ reorders it from $j$-major to $d$-major. Applying it to the rows and columns of $\mathcal{K}$ results in a block matrix $\mathcal{K}^{\prime}\equiv\Pi\mathcal{K}\Pi^{\top}$ that satisfies $\mathcal{K}_{ij,uv}=\mathcal{K}^{\prime}_{uv,ij}$. Section C illustrates this operator with some examples.
Observation 1.
The permutation $\Pi$ is involutory, meaning $\Pi^{-1}=\Pi$.
This implies that to solve a system $\mathcal{K}x=b$, we can instead solve $\mathcal{K}^{\prime}x^{\prime}=\Pi b$ for $x^{\prime}$, then report $x=\Pi x^{\prime}$.
This permutation reduces the bandwidth of certain block matrices:
Observation 2.
When the blocks of $\mathcal{K}$ are banded with bandwidth $w$, $\mathcal{K}^{\prime}$ is block-banded with bandwidth $w$.
The blocks in $\mathcal{K}$ are either block-diagonal, block-upper-bi-diagonal, or block-lower-bi-diagonal, so $\mathcal{K}^{\prime}$ is at most block-tri-diagonal. Denote the $ij$th block of $\mathcal{K}^{\prime}$ as $B_{ij}$:
$B_{ij}\equiv\begin{bmatrix}K_{11,ij}&K_{12,ij}&K_{13,ij}\\ K_{21,ij}&K_{22,ij}&K_{23,ij}\\ K_{31,ij}&K_{32,ij}&K_{33,ij}\\ \end{bmatrix}$.
Then $\mathcal{K}^{\prime}$ can be written as the block-tri-diagonal matrix
$\mathcal{K}^{\prime}=\begin{bmatrix}B_{11}&B_{12}&0&\cdots&&&0\\ B_{21}&B_{22}&B_{23}&\ddots&&&\vdots\\ 0&B_{32}&B_{33}&B_{34}&\ddots&&\\ \vdots&\ddots&\ddots&\ddots&\ddots&\ddots&\vdots\\ 0&\cdots&0&0&B_{L-1,L-2}&B_{L-1,L-1}&B_{L-1,L}\\ 0&\cdots&0&0&0&B_{L,L-1}&B_{L,L}\end{bmatrix}$. (14)
That means we can solve the system $\mathcal{K}^{\prime}x^{\prime}=b^{\prime}$ by first decomposing $\mathcal{K}^{\prime}$ via block LDU decomposition into the product of a lower-block-bi-diagonal matrix, a block-diagonal matrix, and an upper-block-bi-diagonal matrix as $LDUx^{\prime}=b^{\prime}$, then solve for $Ly=b^{\prime}$ using back substitution, and finally solve for $DUx^{\prime}=y$ using forward substitution. Because these systems are block-bi-diagonal, solving them looks like forward and backward propagation on a chain. The running time for these solvers is linear in the depth of the network and cubic in the dimension of each block $B_{ij}$. Depending on the block, the dimension of each of these blocks is either in the number of parameters or the number of activations in the corresponding layer. In all,assuming $a$ and $p$ are the largest activation and parameter count for a layer, the running time for solving this system is $O\left(L\max(a,p)^{3}\right)$ operations.
7 Conclusion
We have described a method to compute the product of the inverse Hessian of a deep neural network with a vector. Compared to the naive method, which stores the Hessian and requires computation cubic in the depth of the network, this method does not store the Hessian and requires computation linear in the depth of the network.
In its final stage, the method relies on a forward and backward substitution to solve a block tridiagonal system. This step bears a similarity to Pearlmutter’s method for computing the Hessian-vector product: It can be interpreted as running backpropagation on a modified version of the original network whose gradient is the desired Hessian-inverse-vector product. The downside of this similarity is that solving large linear systems with LDU factorization is prone to numerical instability. An improvement on the proposal is to solve the block tridiagonal system with an off-the-shelf banded solver.
Our hope is to use this technique as a preconditioner to speed up stochastic gradient descent. Since the method’s speedup is greatest when the network is tall and skinny, we hope it might rekindle interest in extremely deep architectures.
Appendix A The Hessian
We use the facts that $dM^{-1}=-M^{-1}(dM)M^{-1}$ and $b=e_{L}M^{-1}$ to write
$dg$ $=g(x+dx)-g(x)$ (15)
$=d(e_{L}M^{-1}D_{x})$
$=e_{L}M^{-1}(dD_{x})+e_{L}\left(dM^{-1}\right)D_{x}$
$=b\cdot dD_{x}-e_{L}M^{-1}(dM)M^{-1}D_{x}$
$=b\cdot dD_{x}-b\cdot(dM)M^{-1}D_{x}$ (16)
We compute each of these terms separately. As part of this agenda, we will rely on the gradient of tensor-valued functions $g:\mathop{\mathbb{R}}^{d}\to\mathop{\mathbb{R}}^{o_{1}\times\cdots\times o_{k}}$. Define this gradient $\nabla_{x}g(x)\in\mathop{\mathbb{R}}^{(o_{1}\cdots o_{k})\times d}$ as the unique matrix-valued function that satisfies
$\mathrm{vec}\left(g(x+dx)-g(x)\right)\to\nabla_{x}g(x)\cdot dx$ (17)
as $dx\to 0$. This convention readily implies the Hessian of a vector-valued function: If $g:\mathop{\mathbb{R}}^{d}\to\mathop{\mathbb{R}}^{o}$, then $\nabla_{xx}g(x)\in\mathop{\mathbb{R}}^{o\times d^{2}}$ is the unique matrix such that $\mathrm{vec}\left(\nabla_{x}g(x+dx)-\nabla_{x}g(x)\right)\to\nabla_{xx}g(x)\;dx$. This convention also readily accommodates the chain rule. For example, the gradient of $h(x)\equiv f(g(x))$ for matrix-valued $f$ and $g$ can be written as $\nabla f\nabla g$ as expected. It also implies partial derivatives like $\nabla_{yz}g$ for $g:\mathop{\mathbb{R}}^{|x|}\to\mathop{\mathbb{R}}^{|g|}$. If $y\in\mathop{\mathbb{R}}^{|y|}$ and $z\in\mathop{\mathbb{R}}^{|z|}$ are restrictions of $x\in\mathop{\mathbb{R}}^{|x|}$ to some $|y|$ and $|z|$-dimensional subsets, then $\nabla_{z}g(x)\in\mathop{\mathbb{R}}^{|g|\times|z|}$, and $\nabla_{yz}g(x)=\nabla_{y}\nabla_{z}g(x)\in\mathop{\mathbb{R}}^{|g||z|\times|y|}$. See Chapter 6 of Magnus & Neudecker [magnus-neudecker] for a deeper treatment of higher order derivatives of vector-valued functions.
A.1 The term involving $dD_{x}$
[…]
$b\cdot(dD_{x})$ $=\begin{bmatrix}b_{1}&\cdots&b_{L}\end{bmatrix}\begin{bmatrix}dD_{1}&&\\ &\ddots&\\ &&dD_{L}\end{bmatrix}$
$=\begin{bmatrix}b_{1}\cdot dD_{1}&\cdots&b_{L}\cdot dD_{L}\end{bmatrix}$
$=\begin{bmatrix}\mathrm{vec}\left(dD_{1}\right)^{\top}\left(I\otimes b_{1}^{\top}\right)&\cdots&\mathrm{vec}\left(dD_{L}\right)^{\top}\left(I\otimes b_{L}^{\top}\right)\end{bmatrix}$
$=\begin{bmatrix}\mathrm{vec}\left(dD_{1}\right)\\ \vdots\\ \mathrm{vec}\left(dD_{L}\right)\end{bmatrix}^{\top}\begin{bmatrix}I\otimes b_{1}^{\top}&&\\ &\ddots&\\ &&I\otimes b_{L}^{\top}\end{bmatrix}$ (18)
Observe that $\mathrm{vec}\left(dD_{\ell}\right)=d\,\mathrm{vec}\nabla_{x}f_{\ell}(z_{\ell-1};x_{\ell})$ varies with $dx$ through both its arguments $x_{\ell}$ and $z_{\ell-1}$. Using mixed partials of vector-valued functions described above, we get
$\mathrm{vec}\left(dD_{\ell}\right)=d\,\mathrm{vec}\left(\nabla_{x}f_{\ell}\right)=\left(\nabla_{xx}f_{\ell}\right)\;dx_{\ell}+\left(\nabla_{zx}f_{\ell}\right)\;dz_{\ell-1}.$ (19)
Stacking these equations gives
$\begin{bmatrix}\mathrm{vec}\left(dD_{1}\right)\\ \vdots\\ \mathrm{vec}\left(dD_{L}\right)\end{bmatrix}$ $=\begin{bmatrix}\nabla_{xx}f_{1}&&\\ &\ddots&\\ &&\nabla_{xx}f_{L}\end{bmatrix}dx+\begin{bmatrix}\nabla_{zx}f_{1}&&\\ &\ddots&\\ &&\nabla_{zx}f_{L}\end{bmatrix}\begin{bmatrix}dz_{0}\\ \vdots\\ dz_{L-1}\end{bmatrix}$. (20)
Each vector $dz_{\ell}$ in turn varies with $dx$ via $dz_{\ell}=(\nabla_{x}f_{\ell})dx_{\ell}+(\nabla_{z}f_{\ell})dz_{\ell-1}$, with the base case $dz_{0}=0$, since the input $z_{0}$ does not vary with $dx$. Stacking up this recurrence gives
$\begin{bmatrix}I&&\\ -\nabla_{z}f_{2}&I&\\ &&\ddots\\ &&-\nabla_{z}f_{L}&1\\ \end{bmatrix}\begin{bmatrix}dz_{1}\\ \vdots\\ dz_{L-1}\\ dz_{L}\end{bmatrix}=\begin{bmatrix}\nabla_{x}f_{1}&&\\ &\ddots&\\ &&\nabla_{x}f_{L}\end{bmatrix}dx.$ (21)
We can solve for the vector $\begin{bmatrix}dz_{1}\\ \vdots\\ dz_{L}\end{bmatrix}=M^{-1}D_{x}dx$ and use the downshifting matrix
$P\equiv\begin{bmatrix}0\\ I&0\\ &\ddots\\ &I&0\end{bmatrix}$ (22)
to plug back the vector $\begin{bmatrix}dz_{0}\\ \vdots\\ dz_{L-1}\end{bmatrix}=PM^{-1}D_{x}dx$:
$\begin{bmatrix}\mathrm{vec}\left(dD_{1}\right)\\ \vdots\\ \mathrm{vec}\left(dD_{L}\right)\end{bmatrix}=\left(\begin{bmatrix}\nabla_{xx}f_{1}&&\\ &\ddots&\\ &&\nabla_{xx}f_{L}\end{bmatrix}+\begin{bmatrix}\nabla_{zx}f_{1}\\ &\ddots&\\ &&\nabla_{zx}f_{L}\end{bmatrix}PM^{-1}D_{x}\right)dx.$ (23)
A.2 The term involving dMdM
The matrix $dM$ is lower-block-diagonal with $dM_{2},\ldots,dM_{L}$, and $dM_{\ell}\equiv d\nabla_{z}f_{\ell}$. Similar to the above, we can write
$b$ $\cdot(dM)M^{-1}D_{x}=\begin{bmatrix}b_{1}&\cdots&b_{L-1}&b_{L}\end{bmatrix}\begin{bmatrix}0\\ -dM_{2}&0\\ &\ddots\\ &-dM_{L}&0\\ \end{bmatrix}M^{-1}D_{x}$ (24)
$=-\begin{bmatrix}b_{2}\cdot dM_{2}&\cdots&b_{L}\cdot dM_{L}&0\end{bmatrix}M^{-1}D_{x}$ (25)
$=-\begin{bmatrix}\mathrm{vec}\left(dM_{2}\right)^{\top}\left(I\otimes b_{2}^{\top}\right)&\cdots&\mathrm{vec}\left(dM_{L}\right)^{\top}\left(I\otimes b_{L}^{\top}\right)&0\end{bmatrix}M^{-1}D_{x}$ (26)
$=-\begin{bmatrix}\mathrm{vec}\left(dM_{1}\right)\\ \vdots\\ \mathrm{vec}\left(dM_{L}\right)\end{bmatrix}^{\top}\begin{bmatrix}0\\ I\otimes b_{2}^{\top}&0\\ &&\ddots\\ &&I\otimes b_{L}^{\top}&0\end{bmatrix}M^{-1}D_{x}$ (27)
$=-\begin{bmatrix}\mathrm{vec}\left(dM_{1}\right)\\ \vdots\\ \mathrm{vec}\left(dM_{L}\right)\end{bmatrix}^{\top}\begin{bmatrix}I\otimes b_{1}^{\top}\\ &\ddots\\ &&I\otimes b_{L}^{\top}\end{bmatrix}PM^{-1}D_{x}.$ (28)
Each matrix $dM_{\ell}=d\nabla_{z}f_{\ell}(z_{\ell-1};x_{\ell})$ varies with $dx$ through both $x_{\ell}$ and $z_{\ell-1}$ as $d\,\mathrm{vec}\left(M_{\ell}\right)=\left(\nabla_{xz}f_{\ell}\right)dx_{\ell}+\left(\nabla_{zz}f_{\ell}\right)dz_{\ell-1}$. Following the steps of the previous section gives
$\begin{bmatrix}\mathrm{vec}\left(dM_{1}\right)\\ \vdots\\ \mathrm{vec}\left(dM_{L}\right)\end{bmatrix}=\left(\begin{bmatrix}\nabla_{xz}f_{1}&&\\ &\ddots&\\ &&\nabla_{xz}f_{L}\end{bmatrix}+\begin{bmatrix}\nabla_{zz}f_{1}\\ &\ddots&\\ &&\nabla_{zz}f_{L}\end{bmatrix}PM^{-1}D_{x}\right)dx.$ (29)
A.3 Putting it all together
We have just shown that the Hessian of the deep net has the form
$H\equiv\frac{\partial^{2}z_{L}}{\partial x^{2}}$ $=D_{D}\left(D_{xx}+D_{zx}PM^{-1}D_{x}\right)+D_{x}^{\top}M^{-T}P^{\top}D_{M}\left(D_{xz}+D_{zz}PM^{-1}D_{x}\right)$ (30)
$=D_{D}D_{xx}+D_{D}D_{zx}PM^{-1}D_{x}+D_{x}^{\top}M^{-\top}P^{\top}D_{M}D_{xz}+D_{x}^{\top}M^{-\top}P^{\top}D_{M}D_{zz}PM^{-1}D_{x}.$ (31)
Appendix B Equation (9) is Pearlmutter’s Hessian-vector multiplication algorithm
We showed that Equation (9) makes it possible to compute Hessian-vector product $Hv$ in time linear in $L$. These operations are equivalent to Pearlmutter’s [pearlmutter94] algorithm, a framework to compute Hessian-vector products in networks with arbitrary topologies. This section specializes Pearlmutter’s machinery to the pipeline topology, and shows that the operations it produces coincide exactly with those of Equation (9).
Consider a set of vectors $v_{1},\ldots,v_{L}$ that match the dimensions of the parameter vectors $x_{1},\ldots,x_{L}$. Just as $z_{L}(x_{1},\ldots,x_{L})$ denotes the loss under the parameters $w$, we’ll consider the perturbed loss $z_{L}(x_{1}+\alpha v_{1},\ldots,x_{L}+\alpha v_{L})$ with a scalar $\alpha$. By the chain rule,
$\frac{\partial}{\partial\alpha}z_{L}(x_{1}+\alpha_{1}v_{1},\ldots,x_{L}+\alpha_{L}v_{L})\bigg|_{\alpha=0}=\nabla_{x}z_{L}(x_{1},\ldots,x_{L})\cdot v.$ (32)
Applying $\nabla_{x}$ to both sides gives
$\nabla_{x}\frac{\partial}{\partial\alpha}z_{L}(x_{1}+\alpha_{1}v_{1},\ldots,x_{L}+\alpha_{L}v_{L})\bigg|_{\alpha=0}=\nabla_{x}^{2}z_{L}(x_{1},\ldots,x_{L})\cdot v.$ (33)
In other words, to compute the Hessian-vector product $\nabla_{x}^{2}z_{L}\cdot v$, it suffices to compute the gradient of $\frac{\partial z_{L}}{\partial\alpha}$ with respect to $x$. We can do this by applying standard backpropagation to $\frac{\partial z_{L}}{\partial\alpha}$. At each stage $\ell$ during its backward pass, backpropagation produces $\frac{\partial}{\partial x_{\ell}}\frac{\partial z_{L}}{\partial\alpha}=\nabla_{x_{\ell},x}z_{L}\cdot v$, yielding a block of rows in $\nabla_{x}^{2}z_{L}\cdot v$.
To see that this generates the same operations as applying Equation (9) to $v$, we’ll write the backprop operations from Equation (2) against $\frac{\partial z_{L}}{\partial\alpha}$ explicitly. We’ll use again the fact that $z_{\ell}$ depends on $\alpha$ through both $z_{\ell-1}$ and $x_{\ell}+\alpha v_{\ell}$ to massage the backward recursion for $\frac{\partial z_{L}}{\partial\alpha}$ into a format that matches Equation (2):
$b^{\prime}_{\ell}$ $\equiv\frac{\partial}{\partial z_{\ell}}\frac{\partial z_{L}}{\partial\alpha}=\frac{\partial}{\partial\alpha}\frac{\partial z_{L}}{\partial z_{\ell}}=\frac{\partial}{\partial\alpha}b_{\ell}=\frac{\partial}{\partial\alpha}\left[b_{\ell+1}\cdot\nabla_{z}f_{\ell+1}\right]$ (34)
$=b^{\prime}_{\ell+1}\cdot\nabla_{z}f_{\ell+1}+\left[\left(I\otimes b_{\ell+1}\right)\frac{\partial}{\partial\alpha}\text{vec}\left(\nabla_{z}f_{\ell+1}\right)\right]^{\top}$ (35)
$=b^{\prime}_{\ell+1}\cdot\nabla_{z}f_{\ell+1}+\left[\left(I\otimes b_{\ell+1}\right)\left(\nabla_{zz}f_{\ell+1}\cdot\frac{\partial z_{\ell}}{\partial\alpha}+\nabla_{xz}f_{\ell+1}\cdot v_{\ell+1}\right)\right]^{\top}.$ (36)
During the backward pass, from $b_{\ell}$ and $b^{\prime}_{\ell}$, we compute
$\frac{\partial}{\partial x_{\ell}}\frac{\partial z_{L}}{\partial\alpha}$ $=\left(\nabla_{x_{\ell},x}z_{L}\right)\cdot v=\frac{\partial}{\partial\alpha}\frac{\partial z_{L}}{\partial x_{\ell}}=\frac{\partial}{\partial\alpha}\left[\frac{\partial z_{L}}{\partial z_{\ell}}\frac{\partial z_{\ell}}{\partial x_{\ell}}\right]=\frac{\partial}{\partial\alpha}\left[b_{\ell}\cdot\nabla_{x}f_{\ell}\right]$ (37)
$=b^{\prime}_{\ell}\cdot\nabla_{x}f_{\ell}+\left[\left(I\otimes b_{\ell}\right)\frac{\partial}{\partial\alpha}\text{vec}\left(\nabla_{x}f_{\ell}\right)\right]^{\top}$ (38)
$=b^{\prime}_{\ell}\cdot\nabla_{x}f_{\ell}+\left[\left(I\otimes b_{\ell}\right)\left(\nabla_{zx}f_{\ell}\cdot\frac{\partial z_{\ell-1}}{\partial\alpha}+\nabla_{xx}f_{\ell}\cdot v_{\ell}\right)\right]^{\top}.$ (39)
Stacking these backward equations horizontally with $b^{\prime}\equiv\left[\begin{smallmatrix}b^{\prime}_{1}&\cdots&b^{\prime}_{L}\end{smallmatrix}\right]$, $g^{\alpha}_{\ell}\equiv\frac{\partial z_{\ell}}{\partial\alpha}$ and $g^{\alpha}\equiv\left[\begin{smallmatrix}g^{\alpha}_{1}\\ \vdots\\ g^{\alpha}_{L}\end{smallmatrix}\right]$, then transposing, gives
$M^{\top}\left(b^{\prime}\right)^{\top}$ $=PD_{M}\left(D_{zz}g^{\alpha}+D_{xz}v\right)$
$\nabla_{x}^{2}z_{L}\cdot v$ $=D_{x}^{\top}\left(b^{\prime}\right)^{\top}+D_{D}\left(D_{zx}Pg^{\alpha}+D_{xx}v\right).$ (40)
$g^{\alpha}_{\ell}$ can be computed during the forward pass via
$g^{\alpha}_{\ell}\equiv\frac{\partial z_{\ell}}{\partial\alpha}=\nabla_{z}f_{\ell}\cdot\frac{\partial z_{\ell-1}}{\partial\alpha}+\nabla_{x}f_{\ell}\cdot v_{\ell}=\nabla_{z}f_{\ell}\cdot g^{\alpha}_{\ell-1}+\nabla_{x}f_{\ell}\cdot v_{\ell},$ (41)
which when stacked up, gives $Mg^{\alpha}=D_{x}v$. Plugging $g^{\alpha}$ back into Equation (B) and solving for $b^{\prime}$ gives
$\nabla_{x}^{2}z_{L}\cdot v=D_{x}^{\top}M^{-\top}PD_{M}\left(D_{zz}M^{-1}D_{x}v+D_{xz}v\right)+D_{D}\left(D_{zx}PM^{-1}D_{x}v+D_{xx}v\right).$ (42)
This coincides with Equation (9), showing that the two algorithms are equivalent.
Appendix C Examples of Pivoting Matrices to Block-Banded Form
Example 1.
Suppose each block $D_{ij}$ of $\mathcal{K}$ is diagonal. Then the full $4\times 4$ matrix is
$\mathcal{K}=\begin{bmatrix}D_{11}&D_{12}\\ D_{21}&D_{22}\end{bmatrix}=\begin{bmatrix}a&0&c&0\\ 0&b&0&d\\ e&0&g&0\\ 0&f&0&h\end{bmatrix}$,
where
$D_{11}=\begin{bmatrix}a&0\\ 0&b\end{bmatrix},\quad D_{12}=\begin{bmatrix}c&0\\ 0&d\end{bmatrix},\quad D_{21}=\begin{bmatrix}e&0\\ 0&f\end{bmatrix},\quad D_{22}=\begin{bmatrix}g&0\\ 0&h\end{bmatrix}$.
The vector $x=\begin{bmatrix}x_{1}^{1}\\ x_{1}^{2}\\ x_{2}^{1}\\ x_{2}^{2}\end{bmatrix}$ (in $j$-major order) is reordered by $\Pi$ into $\begin{bmatrix}x_{1}^{1}\\ x_{2}^{1}\\ x_{1}^{2}\\ x_{2}^{2}\end{bmatrix}$ ($v$-major order). Then
[…]
Notice that the resulting matrix has become block-diagonal, whereas $\mathcal{K}$ had a bandwidth of 2.
Example 2.
Now endow each block $D_{ij}$ with an upper off-diagonal entry:
$D_{11}=\begin{bmatrix}a&\alpha\\ 0&b\end{bmatrix},\quad D_{12}=\begin{bmatrix}c&\beta\\ 0&d\end{bmatrix},\quad D_{21}=\begin{bmatrix}e&\delta\\ 0&f\end{bmatrix},\quad D_{22}=\begin{bmatrix}g&\gamma\\ 0&h\end{bmatrix}$.
The full $4\times 4$ matrix is
$\mathcal{K}=\begin{bmatrix}D_{11}&D_{12}\\ D_{21}&D_{22}\end{bmatrix}=\begin{bmatrix}a&\alpha&c&\beta\\ 0&b&0&d\\ e&\delta&g&\gamma\\ 0&f&0&h\end{bmatrix}$.
After applying the same permutation $\Pi$ as before, we obtain
$\Pi\mathcal{K}\Pi^{\top}=\begin{bmatrix}a&c&\alpha&\beta\\ e&g&\delta&\gamma\\ 0&0&b&d\\ 0&0&f&h\end{bmatrix}$.
The off-diagonal entries in each block $D_{ij}$ are collected in the upper-right block of the permuted matrix. The permuted matrix is block-banded with bandwidth 2, with a dense upper-right block and a sparse lower-left block.